The fan blades are rotating with an angular speed of -120 rad/s . As the plane takes off, the angular velocity of the blades reaches -390 rad/s in a time of 19s. Find the angular acceleration, assuming it to be constant.




1
Expert's answer
2022-01-07T09:28:36-0500

Answer

Initial angular speed

"\omega_0=-120rad/sec"

After t=19 sec angular speed

"\omega=-390 rad/sec"

Now newton equation of motion of circular motion

Then angular acceleration

"\alpha=\frac{\omega-\omega_0}{t}"

Putting all values

"\alpha=\frac{-390+120}{19}=14.21 rad/sec^2"